OpenStudy (bahrom7893):

write e^(-2+i) in a+bi form

OpenStudy (bahrom7893):

Is it: (1/e^2)*(Cos1 + iSin1) ?

OpenStudy (bahrom7893):

@satellite73

OpenStudy (bahrom7893):

a complex number in form: z = r*e^(i*theta) = r(Cos(theta)+iSin(theta))

OpenStudy (bahrom7893):

@phi

OpenStudy (swissgirl):

\( e^{i\theta} = Cos\theta + iSin\theta\) Ya thats what i Have seen around so I am assuming ur answer is correct

OpenStudy (bahrom7893):

I guess.. it's just that I thought it would look somewhat neater

OpenStudy (bahrom7893):

well thanks everyone

OpenStudy (bahrom7893):

back to studying..

OpenStudy (anonymous):

start with \[e^{-2}e^i\] then oh, you get what you wrote

OpenStudy (anonymous):

since \(e^{ix}=\cos(x)+i\sin(x)\)

OpenStudy (kc_kennylau):

I've recently dealt with this kinda things before... let me think...

OpenStudy (kc_kennylau):

Wait oh satellite has already solved the problem, a=0 and b=\(\large e^{-2}\)